Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
takafumi.ushiba - 23:20 Monday 31 October 2022 (22749) Print this report
PRMI lock trial Day3: PRMI lock trial with 1f signals at POP and ADS enggement with PRMI

Continued work of klog22726.

I tried to lock PRMI with 1f signals at POP tables (POP17Q for MICH and POP45I for PRCL).
POP90 bildup was observed and I tried to engage ADS for maximize it but not yet finished

As usual, after engineers alignmed the mirrors for FPMI, I aligned the PRM and IMMT2 to maximize POP90 signals when PRX was locked with POP45I signals.
After that, I started to lock trial of FPMI with 1f signals.
Since the optical gains are drastically changed when the sideband was resonating PRC, I used POP90 for normalize the RFPD signals, which is same as O3GK era.

After several trial (I have no idea on systematic way to decide the filter gain, I just tried several settings), Icould lock FPMI with 1f signals.
Then, I tried to engage ADS for IMMT2 and PRM to maximize POP90I, and for IMMT1 to maximize POPDC.
Figure 1shows the time series during engagement of ADSs.
Sideband bildup (left bottom: POP90I) are increased from 0.2 to 0.6 by ADSs.

However, after a while, POP90 signals are decreased and lock was lost.
I have no time to confirm the reason but maybe some of the ADSs are not good in the current gain setup.
I will investigate it tomorrow.

Figure 2 shows the today's ADS setting not to forget the current status.

Images attached to this report
Search Help